QUANTUM CHEMISTRY AND COMPUTATION

\underline{WILLIAM H KIRCHHOFF}, NIST, 100 Bureau Drive, Stop 8370, Gaithersburg, MD 20899-8370.

The springboard for this talk will be the oft-quoted statement by the mathematician Richard Hamming, ”The purpose of computing is insight, not numbers.” Beginning with Pitzer’s thesis on the barrier to internal rotation in ethane and continuing on to the present state of computational chemistry, the question of what we have learned and whether or not we like what we have learned will be discussed.